OpenStudy (anonymous):

Simplify the given expression. square root of negative 10 end root times square root of negative 8

OpenStudy (anonymous):

\[\sqrt{-10} \times \sqrt{-8}\]

OpenStudy (anonymous):

options??

OpenStudy (anonymous):

let me write them

OpenStudy (anonymous):

\[-4\sqrt{5}\] \[4\sqrt{5}\] \[4i \sqrt{5}\] \[-4i \sqrt{5}\]

OpenStudy (anonymous):

\[i*\sqrt{10}*\sqrt{8}*i\] \[-\sqrt{80}=-4\sqrt{5}\]

hero (hero):

@mscrosscountry , did you understand all of @Sarkar steps?